# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@
Abstract.
We study Furstenberg set problem, and exceptional set estimate for Marstrand’s orthogonal projection in prime fields for all dimensions. We define the Furstenberg index $\mathbf{F}(s,t;n,k)$ and the Marstrand index $\mathbf{M}(a,s;n,k)$.
It is shown that the two-dimensional result for Furstenberg set problem implies all higher dimensional results.
Key words and phrases:
Furstenberg problem
2020 Mathematics Subject Classification:
28A75, 28A78
1. Introduction
We study the Furstenberg set problem, as well as the exceptional set estimate in prime fields. The main novelty of this paper is that we prove, for these two problems, the two-dimensional result implies all higher dimensional results. In common sense, high dimensional version is harder than low dimensional version. For example, a closely related problem—Kakeya conjecture is known for dimension two but wide open for dimension bigger than two. Hence, it is surprising that for these two problems in prime fields, all higher dimensional results can be deduced from the two dimensional result.
1.1. Background
We first introduce the two problems—Furstenberg set problem and exceptional set estimate in Euclidean space.
1.1.1. Furstenberg set problem
The Furstenberg set problem was originally considered by Wolff [16] in $\mathbb{R}^{2}$. Fix $s\in(0,1],t\in(0,2]$. We say $E\subset\mathbb{R}^{2}$ is a $(s,t)$-set in $\mathbb{R}^{2}$, if $E$ is of the form
@@ 1.1.1. Furstenberg set problem
where $\mathcal{L}\subset A(1,2)$ is a set of lines with $\dim\mathcal{L}=t$ and for each $\ell\in\mathcal{L}$ there exists $Y(\ell)\subset\ell$ with $\dim Y(\ell)=s$. ($\dim$ will always mean Hausdorff dimension.) Furstenberg set problem asks about the optimal lower bound on the Hausdorff dimension of $(s,t)$-sets, i.e.,
[…]
@@
[…]
For more reference and the history on this problem, we refer to [15], [13].
Furstenberg set problem was also considered in higher dimensions, see for example [5], [1], [7]. It can be formulated similarly as follows.
@@
where $\mathcal{V}\subset A(k,n)$ is a set of $k$-planes with $\dim\mathcal{V}=t$ and for each $V\in\mathcal{V}$ there exists $Y(V)\subset V$ with $\dim Y(V)=s$. Furstenberg set problem in $\mathbb{R}^{n}$ for $k$-planes asks to compute
[…]
The Furstenberg set problems in higher dimensions remains wide open.
1.1.2. Exceptional set estimate
Next, we introduce the problem of exceptional set estimate. We first introduce this problem over $\mathbb{R}$. The projection theory dates back to Marstrand [8], who showed that if $A$ is a Borel set in $\mathbb{R}^{2}$, then the projection of $A$ onto almost every one-dimensional subspace has Hausdorff dimension $\min\{1,\dim A\}$. This was generalized to $\mathbb{R}^{n}$ by Mattila [9] who showed that if $A$ is a Borel set in $\mathbb{R}^{n}$, then the projection of $A$ onto almost every $k$-dimensional subspace has Hausdorff dimension $\min\{k,\dim A\}$.
One can consider a refined version of the Marstrand’s projection problem, which is known as the exceptional set estimate.
@@ 1.1.2. Exceptional set estimate
The exceptional set estimate is about finding optimal upper bound of $\dim(E_{s}(A;n,k))$ (in terms of $\dim A,s,n,k$), i.e.,
@@
Many progress have been made on this problem, see for example [6], [2], [14], [12], [13].
In $\mathbb{R}^{2}$, D. Oberlin [10] conjectured that for $0<s\leq\min\{1,a\}$ and $A\subset\mathbb{R}^{2}$ with $\dim A=a$, one has the following exceptional set estimate
(2) $\dim(E_{s}(A;2,1))\leq\max\{2s-a,0\}.$
This conjecture was proved by Ren and Wang in [15].
1.2. The problems in prime fields
In the rest of the paper, we only look at prime fields. Let $p$ be a prime and $\mathbb{F}_{p}$ be the prime field with $p$ elements. Throughout the paper, $A\lessapprox B$ means $A\leq C(\log p)^{C}B$ for some large constant $C$ (independent of $p$). $A\approx B$ means both $A\lessapprox B$ and $B\lessapprox A$. $A\lesssim B$ means $A\leq CB$ for a large constant $C$ (independent of $p$). $A\sim B$ means both $A\lesssim B$ and $B\lesssim A$.
1.2.1. Furstenberg set problem
Definition 1 ($(s,t;n,k)$-set).
@@ 1.2.1. Furstenberg set problem
Here, $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ is a set of $k$-planes with #$\#\mathcal{V}\geq\lambda p^{t}$; for each $V\in\mathcal{V}$, $Y(V)$ is a subset of $V$ with #$\#Y(V)\geq\lambda p^{s}$.
Remark 2.
We remark that $s,t,n,k$ are important parameters, while $\lambda$ is not. Since $\lambda$ is not an important parameter, from now on, we will drop $\lambda$ from notation and simply call $(s,t;n,k;\lambda)$-set as $(s,t;n,k)$-set, thinking of $\lambda$ as a constant in $(0,1]$.
We make the following definition.
Definition 3.
We say $(s,t;n,k)$ is admissible, if $1\leq k<n$ are integers and $0\leq s\leq k$, $0\leq t\leq(k+1)(n-k)$.
The Furstenberg set problem can be formulated as:
@@
(3) $\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon}$
@@
The $\gtrsim_{\varepsilon}$ in the above inequality is actually $\gtrsim_{s,t,n,k,\varepsilon}$. We just view $s,t,n,k$ as fixed parameters, so we simplify it to $\lesssim_{\varepsilon}$. The crucial point is that the implicit constant in $\gtrsim$ or $\lesssim$ throughout this paper never depends on $p$, but can depend on any other fixed parameters. We are interested in the behaviour as $p$ goes to infinity, so the goal is to find the index $\mathbf{F}(s,t;n,k)$ for fixed $(s,t;n,k)$.
We see the problem consists of two parts: for fixed $(s,t;n,k)$,
  1. (1)
    (Lower bound) show for any $(s,t;n,k)$-set $E$, we have #$\#E\gtrsim p^{\mathbf{F}(s,t;n,k)-\varepsilon}$;
  2. (2)
    (Upper bound) show the existence of a $(s,t;n,k)$-set $E$, such that #$\#E\lesssim p^{\mathbf{F}(s,t;n,k)}$.
Based on the Ren-Wang’s result (1), it is reasonable to make the conjecture in $\mathbb{F}_{p}^{2}$.
Conjecture 4.
@@
#$\#E\gtrsim_{\varepsilon}p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}-\varepsilon},$
@@
Remark 5.
It is crucial to use prime field $\mathbb{F}_{p}$. Otherwise, the conjecture fails. Consider the problem in $\mathbb{F}_{p}^{2}$, where $q=p^{2}$ is the square of a prime. We choose the set of lines $\mathcal{V}=\{y=ax+b:a,b\in\mathbb{F}_{p}\}$, so #$\#\mathcal{V}=q$ and hence $t=1$. For each line $V:y=ax+b$ in $\mathcal{V}$, define $Y(V):=\{(x,ax+b):x\in\mathbb{F}_{p}\}$, so #$\#Y(V)=p=q^{\frac{1}{2}}$ and hence $s=\frac{1}{2}$. One can calculate #$\#\big{(}\bigcup_{V\in\mathcal{V}}Y(V)\big{)}=p^{2}=q$. However, $\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}=\frac{5}{4}$.
Usually, people believe the problem in finite field is easier than in Euclidean space. However, Conjecture 4 remains open, while Euclidean version have been resolved.
In the paper, we explicitly find the function $\mathbf{F}(s,t;n,k)$. The definition for $\mathbf{F}(s,t;n,k)$ is tricky, so we postpone it later in Definition 11. The following are main results.
Theorem 6 (Upper bound).
@@
#$\#E\lesssim p^{\mathbf{F}(s,t;n,k)}.$
Theorem 7 (Lower bound).
Suppose $(s,t;n,k)$ is admissible. If we assume Conjecture 4 holds, then for any $(s,t;n,k)$-set $E$ over $\mathbb{F}_{p}$, we have
#$\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$
@@
1.2.2. Exceptional set estimate
We also study the exceptional set estimate in prime fields in all dimensions. We show that the two-dimensional result for Furstenberg set problem implies all higher dimensional results for exceptional set estimates.
To define the projection in $\mathbb{F}_{p}^{n}$, we first consider an equivalent definition of projection in $\mathbb{R}^{n}$.
Recall $\pi_{V}:\mathbb{R}^{n}\rightarrow V$ is the orthogonal projection. For $V\in G(k,\mathbb{R}^{n})$, we also define the map
@@ 1.2.2. Exceptional set estimate
as $\pi^{*}_{V}(x):=x+V$. We see that $\pi_{V}^{*}(x)$ is the unique $k$-plane in $A(k,\mathbb{R}^{n})$ that is parallel to $V$ and passes through $x$. It is not hard to see that $\pi_{V}$ and $\pi^{*}_{V^{\perp}}$ are the same things. This motivates the definition of projections in $\mathbb{F}_{p}^{n}$. We will still use the notations $E_{s}(A;n,k),\pi_{V}^{*}$ for $\mathbb{F}_{p}^{n}$ (which have been used for $\mathbb{R}^{n}$ in previous subsection). This will not cause any confusion.
Definition 8 (Exceptional set).
@@
We explicitly find the function $\mathbf{M}(a,s;n,k)$, whose definition is tricky and therefore postponed to Definition 14. We state the following results.
Theorem 9 (Lower bound).
Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{n}$ with #$\#A\gtrsim p^{a}$, such that
(4) #$\#E_{s}(A;n,k)\gtrsim p^{\mathbf{M}(a,s;n,k)}.$
Theorem 10 (Upper bound).
Assume Conjecture 4 holds. Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. If $p$ is large enough, then for any $A\subset\mathbb{F}_{p}^{n}$ with #$\#A\gtrsim p^{a}$, we have
(5) #$\#E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n,k)},$
@@
1.3. Strategy of the proof
The proof of Theorem 6 and Theorem 9 is by constructing examples. The proof of Theorem 7 and Theorem 10 is by a sequence of pigeonholing arguments and induction on both $n$ and $k$. We explain the idea of the proof of Theorem 7.
We consider the case $(n,k)=(3,1)$. Suppose for any $(s^{\prime},t^{\prime};2,1)$-set $E^{\prime}$, we have #$\#E^{\prime}\gtrsim p^{\mathbf{F}(s^{\prime},t^{\prime};2,1)}$. The goal is to prove for any $(s,t;3,1)$-set $E$, #$\#E\gtrsim p^{\mathbf{F}(s,t;3,1)}$. We will carefully analyze the structure of $E$ and find subsets of $E$ which are $(s^{\prime},t^{\prime};2,1)$-set. Then we use the estimate for those $(s^{\prime},t^{\prime};2,1)$-sets.
Write $E=\bigcup_{L\in\mathcal{L}}Y(L)$, where $\mathcal{L}\subset A(1,\mathbb{F}_{p}^{3})$ with #$\#\mathcal{L}=p^{t}$ and $Y(L)\subset L$ with #$\#Y(L)=p^{s}$. Project $\mathcal{L}$ to $\mathbb{F}_{p}^{2}$, and partition $\mathcal{L}$ according to their images. For each $W\in A(1,\mathbb{F}_{p}^{2})$, define $\mathcal{L}_{W}$ to be the set of those $L\in\mathcal{L}$ whose image under the projection $\mathbb{F}_{p}^{3}\rightarrow\mathbb{F}_{p}^{2}$ is $W$. We may assume all the lines in $\mathcal{L}$ are not parallel to $(0,0,1)$ (by throwing away a small portion of $\mathcal{L}$). We obtain the partition

### v2

# Exceptional set estimate in prime fields: dimension two implies higher dimensions

@@
Abstract
In this paper, we study in prime fields the exceptional set estimates, which can be viewed as a refinement of Marstrand’s orthogonal projection theorem. Additionally, we address a Furstenberg-type problem, which is closely related. It is shown that the two-dimensional result implies all higher-dimensional results in the prime-field setting.
[…]
[classification=text]
1 Introduction
We study the exceptional set estimate and a Furstenberg-type problem in prime fields. The main novelty of this paper is that we prove, for these two problems, that the two-dimensional result implies all higher-dimensional results.
1.1 Background
We first introduce the two problems—the Furstenberg set problem and the exceptional set estimate in Euclidean space.
1.1.1 Furstenberg set problem
The Furstenberg set problem was originally considered by Wolff [20] in $\mathbb{R}^{2}$. Fix $s\in(0,1],t\in(0,2]$. We say $E\subset\mathbb{R}^{2}$ is an $(s,t)$-set in $\mathbb{R}^{2}$, if $E$ is of the form
@@ 1.1.1 Furstenberg set problem
where $\mathcal{L}\subset A(1,\mathbb{R}^{2})$ is a set of lines in $\mathbb{R}^{2}$ with $\dim\mathcal{L}=t$ and for each $\ell\in\mathcal{L}$ there exist $Y(\ell)\subset\ell$ with $\dim Y(\ell)=s$. ($\dim$ will always mean Hausdorff dimension.) The Furstenberg set problem asks about the optimal lower bound on the Hausdorff dimension of $(s,t)$-sets, that is,
$\inf_{E\textup{~is~a~}(s,t)\textup{-set}}\dim(E).$
@@
$\inf_{E\textup{~is~a~}(s,t)\textup{-set}}\dim(E)=\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}.$ (1)
For more references and history on this problem, we refer to [19], [17].
The Furstenberg set problem was also considered in higher dimensions; see, for example, [9], [1], [11]. It can be formulated similarly as follows.
@@
where $\mathcal{V}\subset A(k,\mathbb{R}^{n})$ is a set of $k$-planes $\mathbb{R}^{n}$ with $\dim\mathcal{V}=t$ and for each $V\in\mathcal{V}$ there exists $Y(V)\subset V$ with $\dim Y(V)=s$. The Furstenberg set problem in $\mathbb{R}^{n}$ for $k$-planes asks us to compute
$\inf_{E\textup{~is~a~}(s,t;n,k)\textup{-set}}\dim(E).$
The Furstenberg set problem in higher dimensions remains wide open.
Remark 1.1.
The original Furstenberg set problem requires the $k$-planes in $\mathcal{V}$ to have distinct directions. We will not address this version, as it is more challenging than Kakeya conjecture which is open in higher dimensions. Though we still use the term “Furstenberg set problem” to refer to this non-direction-distinct version addressed in this paper, and we hope this does not cause confusion. For progress on the direction-distinct Furstenberg set problem in finite fields, see [2], [3], [4], [5].
1.1.2 Exceptional set estimate
Next, we introduce the problem of the exceptional set estimate. We first introduce this problem over $\mathbb{R}$. The projection theory dates back to Marstrand [12], who showed that if $A$ is a Borel set in $\mathbb{R}^{2}$, then the projection of $A$ onto almost every one-dimensional subspace has Hausdorff dimension $\min\{1,\dim A\}$. This was generalized to $\mathbb{R}^{n}$ by Mattila [13] who showed that if $A$ is a Borel set in $\mathbb{R}^{n}$, then the projection of $A$ onto almost every $k$-dimensional subspace has Hausdorff dimension $\min\{k,\dim A\}$.
One can consider a refined version of Marstrand’s projection problem, which is known as the exceptional set estimate.
@@ 1.1.2 Exceptional set estimate
The exceptional set estimate is about finding the optimal upper bound of $\dim(E_{s}(A;n,k))$ (in terms of $\dim A,s,n,k$), that is,
@@
For technical reasons, $A$ is always assumed to be Borel.
Much progress has been made on this problem; see, for example, [10], [6], [18], [16], [17].
In $\mathbb{R}^{2}$, D. Oberlin [14] conjectured that for $0<s\leq\min\{1,a\}$ and Borel set $A\subset\mathbb{R}^{2}$ with $\dim A=a$, one has the following exceptional set estimate
$\dim(E_{s}(A;2,1))\leq\max\{2s-a,0\}.$ (2)
This conjecture was proved by Ren and Wang in [19] as a corollary of (1).
1.2 The problems in prime fields
In the remainder of the paper, we look only at the prime fields. Let $p$ be a prime and $\mathbb{F}_{p}$ be the prime field with $p$ elements. Throughout the paper, $A\lessapprox B$ means $A\leq C(\log p)^{C}B$ for some large constant $C$ (independent of $p$). $A\approx B$ means both $A\lessapprox B$ and $B\lessapprox A$. $A\lesssim B$ means $A\leq CB$ for a large constant $C$ (independent of $p$). $A\sim B$ means both $A\lesssim B$ and $B\lesssim A$.
1.2.1 Furstenberg set problem
Definition 1.2 $($s,t;n,k)(s,t;n,k)-set).
@@ 1.2.1 Furstenberg set problem
Here, $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{n})$ is a set of $k$-planes with $\mathcal{V}\geq\lambda p^{t}$; for each $V\in\mathcal{V}$, $Y(V)$ is a subset of $V$ with $Y(V)\geq\lambda p^{s}$.
Remark 1.3.
We remark that $s,t,n,k$ are important parameters, while $\lambda$ is not. Since $\lambda$ is not an important parameter, from now on we will drop $\lambda$ from the notation and simply call $(s,t;n,k;\lambda)$-set as $(s,t;n,k)$-set, thinking of $\lambda$ as a fixed constant in $(0,1]$.
We give the following definition.
Definition 1.4.
We say $(s,t;n,k)$ is admissible if $1\leq k<n$ are integers and $0\leq s\leq k$, $0\leq t\leq(k+1)(n-k)$.
The Furstenberg set problem can be formulated as follows:
@@
$\inf_{E:\ (s,t;n,k)\textup{-set}}\#E\gtrsim_{s,t,n,k,\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon}$ (3)
@@
In the above inequality, the implicit constant may depend on many parameters, but never depends on $p$. For simplicity, we may view $s,t,n,k$ as fixed parameters and write $\gtrsim_{s,t,n,k,\varepsilon}$ as $\gtrsim_{\varepsilon}$ throughout the paper. The dependence on $s,t,n,k$ will be mentioned when needed. We are interested in the behavior as $p$ goes to infinity, so the goal is to find the index $\mathbf{F}(s,t;n,k)$ for fixed $(s,t;n,k)$.
We see that the problem consists of two parts: for fixed $(s,t;n,k)$,
  1. 1.
    (Lower bound) show for any $(s,t;n,k)$-set $E$, we have $E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon}$;
  2. 2.
    (Upper bound) show the existence of a $(s,t;n,k)$-set $E$, such that $E\lesssim p^{\mathbf{F}(s,t;n,k)}$.
Based on Ren-Wang’s result (1), it is reasonable to make a conjecture in $\mathbb{F}_{p}^{2}$.
Conjecture 1.5.
@@
$E\gtrsim_{s,t,\varepsilon}p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}-\varepsilon},$
@@
Remark 1.6.
It is crucial to use the prime field $\mathbb{F}_{p}$. Otherwise, the conjecture fails. Consider the problem in $\mathbb{F}_{q}^{2}$, where $q=p^{2}$ is the square of a prime. We choose the set of lines $\mathcal{V}=\{y=ax+b:a,b\in\mathbb{F}_{p}\}$, so $\mathcal{V}=q$ and hence $t=1$. For each line $V:y=ax+b$ in $\mathcal{V}$, define $Y(V):=\{(x,ax+b):x\in\mathbb{F}_{p}\}$, so $Y(V)=p=q^{\frac{1}{2}}$ and therefore $s=\frac{1}{2}$. We can calculate $\big(\bigcup_{V\in\mathcal{V}}Y(V)\big)=p^{2}=q$. However, $\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}=\frac{5}{4}$.
Usually, people believe that the problem in a finite field is easier than in Euclidean space. However, Conjecture 1.5 remains open, whereas the Euclidean version has been resolved.
In the paper, we explicitly find the function $\mathbf{F}(s,t;n,k)$. The definition of $\mathbf{F}(s,t;n,k)$ is tricky, so we postpone it later in Definition 2.1. The following are the main results.
Theorem 1.7 (Upper bound).
@@
$E\lesssim p^{\mathbf{F}(s,t;n,k)}.$
Theorem 1.8 (Lower bound).
Suppose $(s,t;n,k)$ is admissible. If we assume Conjecture 1.5 holds, then for any $(s,t;n,k)$-set $E$ over $\mathbb{F}_{p}$, we have
$E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;n,k)-\varepsilon},$
@@
1.2.2 Exceptional set estimate
We also study the exceptional set estimate in prime fields in all dimensions. We show that the two-dimensional result for the Furstenberg set problem implies all higher-dimensional results for exceptional set estimates.
To define the projection in $\mathbb{F}_{p}^{n}$, we first consider an equivalent definition of the projection in $\mathbb{R}^{n}$.
Recall that $\pi_{V}:\mathbb{R}^{n}\rightarrow V$ is the orthogonal projection. For $V\in G(k,\mathbb{R}^{n})$, we also define the map
@@ 1.2.2 Exceptional set estimate
as $\pi^{*}_{V}(x):=x+V$. We see that $\pi_{V}^{*}(x)$ is the unique $k$-plane in $A(k,\mathbb{R}^{n})$ that is parallel to $V$ and passes through $x$. It is not hard to see that $\pi_{V}$ and $\pi^{*}_{V^{\perp}}$ are the same things. This motivates the definition of projections in $\mathbb{F}_{p}^{n}$. We will still use the notations $E_{s}(A;n,k),\pi_{V}^{*}$ for $\mathbb{F}_{p}^{n}$ (which has been used for $\mathbb{R}^{n}$ in the previous subsection). This will not cause confusion.
Definition 1.9 (Exceptional set).
@@
We explicitly find the function $\mathbf{M}(a,s;n,k)$, whose definition is tricky and therefore postponed to Definition 2.4. We state the following results.
Theorem 1.10 (Lower bound).
Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{n}$ with $A\gtrsim p^{a}$, such that
$E_{s}(A;n,k)\gtrsim p^{\mathbf{M}(a,s;n,k)}.$ (4)
Theorem 1.11 (Upper bound).
Assume Conjecture 1.5 holds. Fix $1\leq k<n$, $a\in(0,n]$ and $0<s$. If $p$ is large enough, then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\gtrsim p^{a}$, we have
$E_{s-\varepsilon}(A;n,k)\lesssim_{\varepsilon}p^{\varepsilon+\mathbf{M}(a,s;n,k)},$ (5)
@@
1.3 Strategy of the proof
The proof of Theorem 1.7 and Theorem 1.10 is by constructing examples. The proof of Theorem 1.8 and Theorem 1.11 is based on a sequence of pigeonhole arguments and induction on both $n$ and $k$. We explain the idea of the proof of Theorem 1.8.
We consider the case $(n,k)=(3,1)$. Suppose that for any $(s^{\prime},t^{\prime};2,1)$-set $E^{\prime}$, we have $E^{\prime}\gtrsim p^{\mathbf{F}(s^{\prime},t^{\prime};2,1)}$. The goal is to prove for any $(s,t;3,1)$-set $E$, $E\gtrsim p^{\mathbf{F}(s,t;3,1)}$. We will carefully analyze the structure of $E$ and find subsets of $E$ which are $(s^{\prime},t^{\prime};2,1)$-set. Then we use the estimate for those $(s^{\prime},t^{\prime};2,1)$-sets.
Write $E=\bigcup_{L\in\mathcal{L}}Y(L)$, where $\mathcal{L}\subset A(1,\mathbb{F}_{p}^{3})$ with $\mathcal{L}=p^{t}$ and $Y(L)\subset L$ with $Y(L)=p^{s}$. Project $\mathcal{L}$ to $\mathbb{F}_{p}^{2}$, and partition $\mathcal{L}$ according to their images. For each $W\in A(1,\mathbb{F}_{p}^{2})$, define $\mathcal{L}_{W}$ to be the set of those $L\in\mathcal{L}$ whose image under the projection $\mathbb{F}_{p}^{3}\rightarrow\mathbb{F}_{p}^{2}$ is $W$. We may assume that all lines in $\mathcal{L}$ are not parallel to $(0,0,1)$ (by throwing away a small portion of $\mathcal{L}$). We obtain the partition